A metal sphere weighing 19.48g is added to 20ml of water in a graduated cylinder. If the density of the metal is 4.50g/ml, what will be the new level of water in the graduated cylinder?
1
Expert's answer
2019-09-04T03:21:46-0400

"V=V_w+V_s"

"V_w=20ml"

"V_s=\frac{m}{\rho}=\frac{19.48}{4.50}=4.33ml"

"V=20+4.33=24.33ml"

Answer 24.33ml